# Hook-Lengths, Symplectic/Orthogonal Contents and Amdeberhan’s Conjectures

### v1

@@ 3.3. Hook-symplectic/orthogonal contents formulas
In this subsection, we prove Theorem 1.1 and Corollary 1.2.
Proof of Theorem 1.1: What we need is to consider the specialization $p_{k}=t,\forall k\geq 1$ of equation (3.1). Equivalently, for any positive integer $n_{1}$, setting
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}\frac{n_{1}+c_{sp}(u)}{h(u)}=\prod_{n=1}^{\infty}\Bigg{(}\frac{1-q^{4n-2}}{1-q^{8n-4}}\cdot\frac{(1-q^{8n-4})^{\binom{n_{1}}{2}}(1-q^{4n-1})^{n_{1}}(1-q^{8n})^{\binom{n_{1}}{2}+n_{1}}}{(1-q^{4n-3})^{n_{1}}(1-q^{8n-6})^{\binom{n_{1}}{2}}(1-q^{8n-2})^{\binom{n_{1}+1}{2}}}\Bigg{)}.$
Thus, we indeed proved the Theorem 1.1 when $t$ is a positive integer. This is enough. Actually, one can expand the both sides of above equation in formal variables $q$. Then the coefficients in the two sides could be regarded as polynomials in $n_{1}$ of finite orders. Since we proved the equivalence of these two polynomials evaluated at infinite many positive integers $n_{1}$, they should be equal to each others as polynomials. That is to say, the above equation holds when replacing $n_{1}$ by any complex number $t$. After doing that, the result is equivalent to equation (1.3). $\Box$
Proof of Corollary 1.2: From the definitions (1.1) and (1.2) of symplectic contents and orthogonal contents, an easy combinatorial fact is that
@@
As a consequence, the weight functions in the left hand side of equations (1.3) and (1.4) are related to each others by
@@
which proves this corollary just by noticing
@@ 4.1. Partition function of double symplectic Schur functions
In this subsection, we obtain a formula for the partition function consisting of double symplectic Schur functions as Theorem 4.2.
We first need the following commutation relation.
@@
The operators $\Gamma_{+}(z)$ and $\exp(G^{*})$ commute to each others. For vertex operator $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$, we have
@@
Proof: By taking dual of Lemma 3.4. $\Box$
@@
Use the symmetric variables $(\bm{y})=(y_{i})$ and write $p_{k}$ as $p_{k}(\bm{y})$, the $k-$power sum of variables $(\bm{y})$. The above Theorem 4.2 is equivalent to the following
@@
Proof: Similar to the proof of Corollary 3.2. One just needs to notice equations (3.3) and the following
@@
Proof of Theorem 4.2: We will prove equation (4.3) below, which is equivalent to Theorem 4.2. First, by equation (3.12), we have
@@
By taking the dual of the last term in the right hand side of above equation, and using the equation (3.14), the left hand side of above equation is equal to
@@
The computation of above vacuum expectation value is similar to the proof of Theorem 3.1. First, we replace the variables $(\bm{p})$ and $(\bm{p}^{\prime})$ by symmetric functions $\big{(}\bm{p}(\bm{y})\big{)}$ and $\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}$ respectively. The commutation relations (2.6) and (2.5) can be used to change the orders of $q^{L_{0}}$, $\Gamma_{+}\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}$ and $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$ to obtain
@@
Then we can use equations (4.1) and (3.7) to change the orders of the operators $\Gamma_{\pm}(\cdot)$, $\exp(G^{*})$ and $\exp(G)$ respectively. The result is that, the left hand side of equation (4.3) is equal to
@@
By the similar method, we can use equations (3.8) and (4.2) to deal with the commutation relations between $\Gamma_{\pm}(\cdot)^{-1}$ and $\exp(G),\exp(G^{*})$. The result is
@@
By doing above process again and again, one obtains that, the left hand side of equation (4.3) is equal to
@@
The vacuum expectation value in above equation can be computed in terms of
@@
where we have used the equation (3.14). This Theorem is thus proved. $\Box$
@@
Proof of Theorem 1.3: For any two positive integers $n_{1}$ and $n_{2}$, taking the special evaluation for equation (4.3) at
@@
Since both sides of above equation could be regarded as elements in the ring $\mathbb{C}[\![q]\!][n_{1},n_{2}]$ and we already proved above equation for all $n_{1},n_{2}\in\mathbb{Z}_{>0}$, above equation should still hold when replacing $n_{1},n_{2}$ by any complex numbers $t_{1},t_{2}$. This first equation (1.5) in this theorem is thus proved.
For the second equation (1.6) in this theorem, one just needs to take the change of variables $(t_{1},t_{2})\rightarrow(-t_{1},-t_{2})$ for equation (1.5) and notices that
@@
which is exactly the first result of Conjecture 6.3 (a) in [1]. The second result of Conjecture 6.3 (a) in [1] similarly follows from special case $t_{1}=-t_{2}=t$ of equation (1.6).
Proof of Corollary 1.4: This is a direct corollary of Theorem (1.1) and Theorem 1.3 as indicated by Amdeberhan [1]. By setting $t=0$ in Theorem (1.1) and $t_{1}=t_{2}=0$ in Theorem 1.3, we have
@@
From the infinite produce formula, it is easy to see that, in above equation, only terms labeled by $q^{n},n\in 2\mathbb{Z}_{\geq 0}$ make non-trivial contribution. Since $(-1)^{n/2}=(-1)^{\binom{n}{2}}$ when $n\in 2\mathbb{Z}_{\geq 0}$, this corollary is obtained by taking the coefficient of $q^{n}$ in the two sides of above equation for an arbitrary fixed nonnegative integer $n$. $\Box$
@@
Actually, Sundaram [18, 19] provided the above equation as an equality for formal functions in variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$ for $l(\lambda)\leq n$. We regard it as an identity in the ring $\mathbb{C}[\![q]\!][\bm{p}(\bm{y}),\bm{p}^{\prime}(\bm{y}^{\prime})]$ and it can also be proved by using the method in deriving Theorem 4.2.
@@
Since both sides of above equation could be regarded as elements in the ring $\mathbb{C}[\![q]\!][n_{1},n_{2}]$, and we already proved the equivalence between them for all $(n_{1},n_{2})\in\mathbb{Z}_{>0}^{2}$, the above equation still holds when replacing $n_{1},n_{2}$ by any complex numbers $t_{1},t_{2}$, which proves this theorem. $\Box$
@@ 5. Conflict of interest and data availability statement
Acknowledgements. The author would like to thank Professors Xiaobo Liu, Jian Zhou and Xiangyu Zhou for encouragement. The author is supported by the National Natural Science Foundation of China (No. 12288201) and the China Postdoctoral Science Foundation (No. 2023M743717).

### v2

@@ 3.3. Hook-symplectic/orthogonal contents formulas
In this subsection, we venture on proving the results in Theorem 1.1 and Corollary 1.2.
Proof of Theorem 1.1: What we need is to consider the specialization $p_{k}=t,\forall k\geq 1$ of equation (3.1). Equivalently, for an arbitrary positive integer $n_{1}$, setting
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}\frac{n_{1}+c_{sp}(u)}{h(u)}=\prod_{n=1}^{\infty}\Bigg{(}\frac{1-q^{4n-2}}{1-q^{8n-4}}\cdot\frac{(1-q^{8n-4})^{\binom{n_{1}}{2}}(1-q^{4n-1})^{n_{1}}(1-q^{8n})^{\binom{n_{1}}{2}+n_{1}}}{(1-q^{4n-3})^{n_{1}}(1-q^{8n-6})^{\binom{n_{1}}{2}}(1-q^{8n-2})^{\binom{n_{1}+1}{2}}}\Bigg{)}.$ (3.18)
Thus, we indeed have proved Theorem 1.1 when $t$ is a positive integer. This is enough. Actually, one can expand both sides of the above equation (3.18) in the formal variable $q$. Then the coefficients on the two sides could be regarded as polynomials in $n_{1}$ of finite orders. Since we proved the equivalence of these two polynomials evaluated at infinite many positive integers $n_{1}$, they should be equal to each other as polynomials. That is to say, the above equation holds when replacing $n_{1}$ by any complex number $t$. After doing that, the result is equivalent to equation (1.3). $\Box$
Proof of Corollary 1.2: Starting with the definitions (1.1) and (1.2) of symplectic and orthogonal contents, an easy combinatorial fact arises as
@@
As a consequence, the weight functions on the left hand side of equations (1.3) and (1.4) are related to each other by
@@
which proves the corollary just by noticing
@@ 4.1. Partition function of double symplectic Schur functions
In this subsection, we obtain a formula for the partition function consisting of double symplectic Schur functions as in Theorem 4.2. We first need the following commutation relation.
@@
The operators $\Gamma_{+}(z)$ and $\exp(G^{*})$ commute with each other. For the vertex operator $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$, we have
@@
Proof: Take the dual of Lemma 3.4. $\Box$
@@
Use the symmetric variables $(\bm{y})=(y_{i})$ and write $p_{k}$ as $p_{k}(\bm{y})$, the $k$-th power sum of the variables $(\bm{y})$. Then, Theorem 4.2 is equivalent to the following.
@@
Proof: This is similar to the proof of Corollary 3.2. One just needs to notice equations (3.3) and the following fact
@@
Proof of Theorem 4.2: Observing that equation (4.3) is equivalent to Theorem 4.2, we will prove (4.3). First, by equation (3.12), we have
@@
By taking the dual of the last term from the right hand side of above equation, and using equation (3.14), the left hand side of the above equation is equal to
@@
The computation of the above vacuum expectation value is similar to the proof of Theorem 3.1. First, we replace the variables $(\bm{p})$ and $(\bm{p}^{\prime})$ by symmetric functions $\big{(}\bm{p}(\bm{y})\big{)}$ and $\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}$, respectively. The commutation relations (2.6) and (2.5) can be used to change the orders of $q^{L_{0}}$, $\Gamma_{+}\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}$ and $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$ to obtain
@@
Then we can use equations (4.1) and (3.7) to change the orders of the operators $\Gamma_{\pm}(\cdot)$, $\exp(G^{*})$ and $\exp(G)$. The result is that, the left hand side of equation (4.3) is equal to
@@
By analogy, we can use equations (3.8) and (4.2) to deal with the commutation relations between $\Gamma_{\pm}(\cdot)^{-1}$ and $\exp(G),\exp(G^{*})$. The result is
@@
By repeating the above process again and again, one obtains that, the left hand side of equation (4.3) is equal to
@@
The vacuum expectation value in the above equation can be computed in terms of
@@
where we have used equation (3.14). This completes the proof of the theorem. $\Box$
@@
Proof of Theorem 1.3: For any two positive integers $n_{1}$ and $n_{2}$, taking the special evaluation of equation (4.3) at
@@
Since both sides of the above equation could be regarded as elements in the ring $\mathbb{C}[\![q]\!][n_{1},n_{2}]$ and we already have proved the above equation for all $n_{1},n_{2}\in\mathbb{Z}_{>0}$, it follows that the equation should still hold true when replacing $n_{1},n_{2}$ by arbitrary complex numbers $t_{1},t_{2}$. Hence, the first equation (1.5) of Theorem 1.3 is thus proved.
For the second equation (1.6) in this theorem, it suffices to make the change of variables $(t_{1},t_{2})\rightarrow(-t_{1},-t_{2})$ in equation (1.5) and to notice that
@@
which is exactly the first result in Conjecture 6.3 (a) of [1]. Similarly, the second result of Conjecture 6.3 (a) in [1] follows from the special special case $t_{1}=-t_{2}=t$ of equation (1.6).
Proof of Corollary 1.4: This is a direct consequence of Theorem 1.1 and Theorem 1.3 as indicated by Amdeberhan [1]. By setting $t=0$ in Theorem 1.1 and $t_{1}=t_{2}=0$ in Theorem 1.3, we have
@@
From the infinite produce formula, it is easy to see that, in above equation, only terms labeled by $q^{n},n\in 2\mathbb{Z}_{\geq 0}$ make non-trivial contribution. Since $(-1)^{n/2}=(-1)^{\binom{n}{2}}$ when $n\in 2\mathbb{Z}_{\geq 0}$, this corollary is obtained by taking the coefficient of $q^{n}$ on both sides of the above equation for an arbitrary fixed nonnegative integer $n$. $\Box$
@@
Actually, Sundaram [18, 19] provided the above equation as an equality for formal functions in the variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$ for $l(\lambda)\leq n$. We regard it as an identity in the ring $\mathbb{C}[\![q]\!][\bm{p}(\bm{y}),\bm{p}^{\prime}(\bm{y}^{\prime})]$ and it can also be proved by using the method employed in deriving Theorem 4.2.
@@
Since both sides of the above equation could be viewed as elements in the ring $\mathbb{C}[\![q]\!][n_{1},n_{2}]$, and we already have proved the equality between them for all $(n_{1},n_{2})\in\mathbb{Z}_{>0}^{2}$, the above equation still holds when replacing $n_{1},n_{2}$ by any complex numbers $t_{1},t_{2}$. The proof is complete. $\Box$
@@ 5. Conflict of interest and data availability statement
Acknowledgements. The author would like to thank Professors Amdeberhan, Andrews, and Ballantine for valuable discussions and their interest in this work. The author would also like to express gratitude to Professors Xiaobo Liu, Jian Zhou, and Xiangyu Zhou for their encouragement. The author is supported by the National Natural Science Foundation of China (No. 12288201) and the China Postdoctoral Science Foundation (No. 2023M743717).